OpenStudy (anonymous):

How do I find the inverse of f(x) when the equation is: f(x)=(x+3)^2, x<-3

OpenStudy (anonymous):

The answer is supposed to be f-1(x) = -√x - 3

OpenStudy (anonymous):

You just replace the viable x with y, and then rearrange the equation. So you have:\[y=(x+3)^{2}\] Replace x with y: \[x=(y+3)^{2}\] rearrange the equation. \[\pm \sqrt{x}=y+3\]\[y=\pm \sqrt{x}-3\] Check to see if the equation is an inverse by substitution. If the solution equals x then it is an inverse. Substituting\[y=-\sqrt{x}-3\] into\[f(x)\] So \[f(-\sqrt{x}-3)=(-\sqrt{x}-3+3)^{2}\] y=x so \[y=-\sqrt{x}-3\] is an inverse. Substituting \[y=\sqrt{x}-3\]into \[f(\sqrt{x}-3)=(\sqrt{x}-3+3)^{2}\] y=x So \[f(\sqrt{x}-3) \] is also an inverse

OpenStudy (anonymous):

Interestingly, when you take both inverses, you get another parabola that fails the verticle line test. So it looks like you take one or the other inverses, but not both.